OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below where A = the pounds of Arabica and R = pounds of Robusta. 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds @mathman806 cam you explain this to me

OpenStudy (anonymous):

ok so first thing that you would do is plug 1000 in for A 700*1000+1200R=1,000,000 multiply 700000+1200R=1,000,000 subtract 700000 1200R=300000 divide by 1200 R=250

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

ok hey I can only be on for another 7 minutes but then I'll be back tomorrow at 12:30

OpenStudy (anonymous):

okay As the number of pages in a photo book increases, the price of the book also increases. There is an additional shipping charge of 15%. The price of a book can be modeled by the equation below where, P = the price of the book, 20 is the printing charge, 0.5 is the charge per page, and x = the number of pages. P = (20 + 0.5x) + 0.15(20 + 0.5x) Jennifer wants to purchase a book but only has $62.10 to spend. What is the maximum number of pages she can have in her book? x = ______________________ pages

OpenStudy (anonymous):

ok just like the last one plug 62.10 in for P 62.10=(20+0.5x)+0.15(20+0.5x) multiply the 0.15 by the 20 and 0.5x 62.10=(20+.05x)+3+0.75x Take parenthesis away 62.1=20+0.5x+3+0.75x combine the sections with same variable 62.1=23+1.25x now solve like how I showed you in the last problem because I gtg bye

OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

i dont get the last question whats the answer

OpenStudy (anonymous):

if 267/3=89 then you could just do 88+89+90=267 so it would be 90